The heaviest human of all time weighed 486 kilograms. This is 12 kilograms less than 6 times Juan's weight. How much does Juan weigh?please help and explai thanks so much!
Answer by checkley77(12844)

Let x be Juan's weight.
486=6X-12
6X=486+12
6X=498
X=498/6
X=83 KG. IS JUAN'S WEIGHT.
Proof:
486=6*83-12
486=498-12
486=486